# Robust Fault-Tolerant Control and Agile Trajectory Planning for Modular Aerial Robotic Systems

### v1

@@ Abstract
Modular Aerial Robotic Systems (MARS) consist of multiple drone units that can self-reconfigure to adapt to various mission requirements and fault conditions. However, existing fault-tolerant control methods exhibit significant oscillations during docking and separation, impacting system stability. To address this issue, we propose a novel fault-tolerant control reallocation method that adapts to arbitrary number of modular robots and their assembly formations. The algorithm redistributes the expected collective force and torque required for MARS to individual unit according to their moment arm relative to the center of MARS mass. Furthermore, We propose an agile trajectory planning method for MARS of arbitrary configurations, which is collision-avoiding and dynamically feasible. Our work represents the first comprehensive approach to enable fault-tolerant and collision avoidance flight for MARS. We validate our method through extensive simulations, demonstrating improved fault tolerance, enhanced trajectory tracking accuracy, and greater robustness in cluttered environments. The videos and source code of this work are available at https://github.com/RuiHuangNUS/MARS-FTCC/
@@ I Introduction
Modular Aerial Robotic Systems (MARS) [1] consist of flying units capable of freely maneuvering in three-dimensional space. They can reconfigure the assembly formation in-flight to facilitate navigating through cluttered and confined environments [2, 3, 4, 5]. Despite the promising applications, agile trajectory planning and control of MARS is rarely explored in the literature that taps into such flexibility. Simply applications of existing generic planning and control algorithms typically lead to poor tracking performances. For example, MARS under a simple PID controller cannot accurately track the collision-free trajectory generated by previous work [6] (See Fig. 1(a)).
@@
Besides, aerial robots are vulnerable to various faults, including multiple propeller failures that jeopardize the control performance and flight safety. This can severely degrade the control performance of MARS. Achieving fault tolerance in various robotic swarms is a challenging problem that has garnered significant attention in recent research [7, 8, 9]. Existing works on MARS primarily focus on docking and separation [10, 11, 12], coordinated flight[9], and self-reconfiguration[8, 13]. However, the shape of MARS may not be well-suited for navigating through narrow urban environments, which limits its adaptability in load-carrying and shape-transforming transportation tasks. Therefore, a method capable of planning arbitrary MARS configurations is required to enhance its versatility and operational efficiency.
@@
Figure 1: MARS is tasked with tracking a collision-free trajectory with one faulty unit. The faulty propellers are marked in red. (a) MARS cannot accurately follow the planned trajectory using an existing collision-free trajectory generation method [6] under a simple PID control. (b) MARS fails to track the trajectory planned with [6] under our proposed fault-tolerant control. (c) MARS can track the trajectory planned with our proposed dynamics-aware planning method relatively accurately under our proposed fault-tolerant control.
@@
Existing studies [14, 6] on collision-free trajectory generation primarily focus on computing two-dimensional and three-dimensional trajectories for arbitrary shapes in complex environments. However, these methods do not consider asymmetrically distributed control outputs, such as thrusters and motors commonly used in aerial and underwater robotics. Consequently, the control performance varies across different directions, causing deviations between the planned collision-free trajectory and the system’s actual tracking capability. Fig. 1(b) shows that MARS fails to track the trajectory planned with [6] under a fault-tolerant controller (FTC) proposed in this paper, as the former does not account for the dynamics in their planning.
@@
To enhance the in-flight safety of MARS, we propose a novel dynamics-aware collision-free trajectory planning algorithm that incorporates FTC during obstacle avoidance. As a preview of our results, we show in Fig. 1(c) that MARS can track the trajectory planned with our proposed dynamics-aware planning method relatively accurately under our proposed FTC. In comparison to Figure 1(b), it highlights the importance of incorporating the dynamics into the trajectory planning and the control design to achieve robust and agile control of MARS.
@@ II-A Units failure
${\rm s.t.}$ $\Sigma u_{a,i}\cdot\mathbf{P}_{i}=0$ (1b)
@@
where $\boldsymbol{u}_{a}=[F_{1},\cdots,F_{i}]$ represents the control output of the MARS, and $F_{i}$ denotes the force produced by the $i$-th normal unit. The function $\text{Var}(\cdot)$ represents the variance. The unit position in the assembly coordinate frame is given by $\mathbf{P}_{i}=[x_{i},y_{i}]$. The desired force and net torque of the MARS are denoted as $F_{\text{assembly}}$ and $\mathbf{M}^{\pm}_{\text{assembly}}=[M_{x}^{+},M_{x}^{-},M_{y}^{+},M_{y}^{-}]$, respectively. The current torque generated by the thrusters is represented as $\boldsymbol{\tau}^{\pm}=[\tau_{x}^{+},\tau_{x}^{-},\tau_{y}^{+},\tau_{y}^{-}]$, which is computed as:
@@ II-B1 Partial Reallocation
where $k$ is the number of normal units. $\Delta F_{\text{failed}}$ and $\Delta\mathbf{M}_{\text{failed}}$ denote the loss of thrust and torque due to rotor failure, respectively, and can be calculated using the control efficiency matrix. Then, the angular velocity $\boldsymbol{\omega}_{i}$ can be determined according to the adaptive configuration control method without failure [9].
@@
where $\mathbf{P}_{n_{r}}$ denotes the original control allocation matrix of a multirotor with $n_{r}$ propellers, and $\mathbf{E}_{i}=\text{diag}\left([n^{-1},|\mathbf{P}_{i}+\mu|\cdot J_{\text{assembly}}/|\sum_{i}^{n}(\mathbf{P}_{i}+\mu)\cdot J_{\text{unit}}|]\right)\in\mathbb{R}^{4\times 4}$ defines the adaptive configuration control allocation parameters. Here, $J$ represents the moment of inertia, and $\mu$ is a small constant introduced to prevent division by zero.
@@ II-B2 Full Reallocation
${\rm s.t.}$ $\Sigma f_{i,j}\cdot\mathbf{P}_{i,j}=0$ (6b)
@@ III Collision-free Trajectory Planning
11:         if $N_{near}\notin C$ and CollisonCheck($N_{near}$then
@@ III-1 Discrete sequence generation
Figure 3: Comparison of trajectory tracking and average position errors across ten repeated experiments with and without fault-tolerant control. The faulty propellers are marked red.
@@
where $R(\phi)$, $R(\theta)$, and $R(\psi)$ represent the rotation matrices about the $z$-, $y$-, and $x$-axes, respectively. The maximum force contributions for each unit and the entire assembly are defined as: $\boldsymbol{u}_{\text{u},i}^{\text{max}}=\begin{bmatrix}f_{i1}^{\text{max}},&f_{i2}^{\text{max}},&f_{i3}^{\text{max}},&f_{i4}^{\text{max}}\end{bmatrix}^{T},\quad\boldsymbol{u}_{a}^{\text{max}}=\begin{bmatrix}\boldsymbol{u}_{\text{u},1}^{\text{max}},&\cdots,&\boldsymbol{u}_{\text{u},n}^{\text{max}}\end{bmatrix}^{T}\in\mathbb{R}^{4n}$.
@@ III-2 Dynamical trajectory
Figure 4: FTC performance during self-reconstruction in $3\times 1$ and $3\times 3$ plus assemblies. Faulty propellers are highlighted in red, while orange arrows indicate the range of oscillation. (a) $3\times 1$ assembly (comparison method). (b) $3\times 1$ assembly (Ours). (c) $3\times 3$ plus assembly (comparison method). (d) $3\times 3$ plus assembly (Ours). (e) and (f) Yaw angle and IMU-measured accelerations of the faulty unit during self-reconfiguration FTC in $4\times 1$ and $3\times 3$ plus assemblies, respectively.
@@ IV-A1 Complete Failure of One and Two Units
To evaluate the performance of the proposed method, a spiral trajectory tracking experiment is conducted. Fig. 3 presents the results of trajectory tracking along with the average position errors over 10 repeated trials. As shown in Fig. 3(a) and (b), when fault-tolerant control is enabled, the assembly remains capable of effectively tracking the spiral direction. However, the tracking trajectory deviates from the target trajectory when faulty units are located far from the center, resulting in higher fault-induced torque. In contrast, without fault-tolerant control, the assembly quickly diverges from the intended trajectory after take-off and fails to complete the tracking task. These results demonstrate that the proposed fault-tolerant control method effectively redistributes power to handle fault scenarios.
Figure 5: Simulation of collision-free trajectory planning for $4\times 1$ and $3\times 2$ assemblies. The annotated diagram illustrates the configuration of MARS, where faulty units are marked in red. The simulation videos are available at https://github.com/RuiHuangNUS/MARS-FTCC/. (a) Normal assembly (SVSDF) [6]. (b) Normal assembly (Ours). (c) Post-failure assembly with FTC (SVSDF). (d) Post-failure assembly with FTC (Ours).
@@
TABLE I: Trajectory Error Analysis: Rotor Failures in 3$\times$2 Assembly.
@@
IV-A2 Rotor Failure and Performance Degradation
Performance degradation is a common type of malfunction, often caused by propeller or rotor failures. To evaluate the fault-tolerant capability of MARS, we performed trajectory tracking experiments under different rotor failure conditions in a 3$\times$2 configuration, as summarized in Table I. Here, $\eta_{2,i}$ represents the coefficient factor of the $i^{th}$ rotor in unit No. 2. The results demonstrate that fault-tolerant control significantly reduces trajectory errors, achieving performance comparable to the fault-free configuration, thereby enabling successful trajectory tracking. Furthermore, when multiple rotors experience reduced thrust, hybrid FTC effectively enhances flight stability. In contrast, the group without internal FTC exhibited considerable trajectory deviations.
@@ IV-C Collision-free Path Planning
Figure 6: Comparison of dynamical trajectory generated by SVSDF [6] and our method. (a) 3$\times$2 assembly (Ours). (b) 3$\times$2 assembly (SVSDF). (c) 4$\times$1 assembly (SVSDF). (d) 4$\times$1 assembly (Ours).
To better demonstrate the advantages of the optimal attitude in simulation environments with dynamic models, we carry out experiments using SVSDF[6] and our method in Coppeliasim[15]. The generated MARS trajectories of the 4$\times$1 and 3$\times$2 assemblies are shown in Fig  6. The attitude-optimized dynamical trajectories minimize abrupt attitude changes and enhance the feasibility of tracking tasks.
The simulation sequence graphs of different methods are shown in Fig. 5. FTC enables MARS to track the planned trajectory after the failure of a unit in both methods, as illustrated in Fig. 5 (c) and (d). However, the trajectory generated by SVSDF poses greater challenges for MARS with dynamic models, often resulting in more pronounced trajectory errors and a higher frequency of collisions, as shown in Fig. 5 (a) and (c). Specifically, in the first row of Fig. 5, a failure occurs in Unit No. 3 of the $4\times 1$ assembly. Our method demonstrates lower tracking errors compared to [6]. When a failure occurs in Unit No. 4, the increased fault torque reduces the assembly’s control capability, leading to higher tracking errors and even collisions, as shown in Fig. 5 (c) for the comparison method [6] in the second row. Furthermore, for the $3\times 2$ assembly, MARS without fault struggles to accurately follow the SVSDF-generated trajectory, as seen in Fig. 5 (a). The performance further deteriorates in the presence of faults, as illustrated in Fig. 5 (c) in the third row. In contrast, our approach, shown in Fig. 5 (b) and (d) in the third row, effectively mitigates collision risks and reduces average trajectory errors, achieving performance comparable to the nominal configuration. These results demonstrate the superiority of our proposed method in ensuring dynamic adaptability and collision-free trajectory planning.
@@ V Conclusion
To achieve robust and safe post-failure control of MARS, this paper proposes a partial and full fault-tolerant power allocation method that redistributes the expected force and torque across all units based on moment arm magnitudes. Compared to existing approaches, our method significantly reduces oscillations during the reconfiguration process. Furthermore, we develop a dynamics-aware, collision-free trajectory planning strategy by extending the A* algorithm to compute optimal attitude nodes that maximize control performance. We then generate a dynamic trajectory incorporating a dynamic feasibility term. Experimental results demonstrate that our approach effectively reduces tracking errors while preventing collisions when following a collision-free trajectory, even in the presence of failures.

### v2

# MARS-FTCP: Robust Fault-Tolerant Control and Agile Trajectory Planning for Modular Aerial Robot Systems

@@ Abstract
Modular Aerial Robot Systems (MARS) consist of multiple drone units that can self-reconfigure to adapt to various mission requirements and fault conditions. However, existing fault-tolerant control methods exhibit significant oscillations during docking and separation, impacting system stability. To address this issue, we propose a novel fault-tolerant control reallocation method that adapts to an arbitrary number of modular robots and their assembly formations. The algorithm redistributes the expected collective force and torque required for MARS to individual units according to their moment arm relative to the center of MARS mass. Furthermore, we propose an agile trajectory planning method for MARS of arbitrary configurations, which is collision-avoiding and dynamically feasible. Our work represents the first comprehensive approach to enable fault-tolerant and collision avoidance flight for MARS. We validate our method through extensive simulations, demonstrating improved fault tolerance, enhanced trajectory tracking accuracy, and greater robustness in cluttered environments. The videos and source code of this work are available at https://github.com/RuiHuangNUS/MARS-FTCP/
@@ I Introduction
Modular Aerial Robot Systems (MARS) [1, 2] consist of flying units capable of freely maneuvering in three-dimensional space. They can reconfigure the assembly formation in-flight to facilitate navigating through cluttered and confined environments [3, 4, 5]. Despite the promising applications, agile trajectory planning and control of MARS is rarely explored in the literature that taps into such flexibility. Simple applications of existing generic planning and control algorithms typically lead to poor tracking performances. For example, MARS under a simple PID controller cannot accurately track the collision-free trajectory generated by previous work [6] (See Fig. 1(a)).
@@
Besides, aerial robots are vulnerable to various faults, including multiple propeller failures that jeopardize the control performance and flight safety [7, 8]. This can severely degrade the control performance of MARS. Achieving fault tolerance in various robotic swarms is a challenging problem that has garnered significant attention in recent research [9, 10, 11]. Existing works on MARS primarily focus on docking and separation [12, 13, 14], coordinated flight[11], and self-reconfiguration[10, 15]. However, the shape of MARS may not be well-suited for navigating through narrow urban environments, which limits its adaptability in load-carrying and shape-transforming transportation tasks. To enhance its versatility and operational efficiency, a method capable of planning arbitrary MARS configurations is required.
@@
Figure 1: MARS is tasked with tracking a collision-free trajectory with one faulty unit. The faulty propellers are marked in red. (a) MARS cannot accurately follow the planned trajectory using an existing collision-free trajectory generation method [6] under a simple PID control and control allocation approach used in previous work [11]. (b) MARS can track the planned trajectory more accurately using our dynamics-aware planning method. (c) MARS fails to track the trajectory planned with [6] under our proposed fault-tolerant control. (d) MARS can track the trajectory planned with our proposed method relatively accurately under our proposed fault-tolerant control.
@@
Existing studies [16, 6] on collision-free trajectory generation primarily focus on computing two-dimensional and three-dimensional trajectories for arbitrary shapes in complex environments. However, these methods do not consider asymmetrically distributed control outputs, such as thrusters and motors commonly used in aerial and underwater robotics. Consequently, the control performance varies across different directions, causing deviations between the planned collision-free trajectory and the system’s actual tracking capability. Fig. 1(c) shows that MARS fails to track the trajectory planned with [6] under a fault-tolerant controller (FTC) proposed in this paper, as the former does not account for the dynamics in their planning.
@@
To enhance the in-flight safety of MARS, we propose a novel dynamics-aware collision-free trajectory planning algorithm that incorporates FTC during obstacle avoidance. As a preview of our results, Fig. 1(b) demonstrates that MARS can track the planned trajectory more accurately using our proposed dynamics-aware planning method, and Fig. 1(d) shows the performance under our proposed FTC scheme. Compared to Fig. 1(a) and (c), these results highlight the importance of incorporating the dynamics into the trajectory planning and the control design to achieve robust and agile control of MARS.
@@ II-A Units failure
${\rm s.t.}\$$\Sigma u_{a,i}\cdot\mathbf{P}_{i}=0$ (1b)
@@
where $\boldsymbol{u}_{a}=[F_{1},\cdots,F_{i}]$ represents the control output of the MARS, and $F_{i}$ denotes the force produced by the $i$-th normal unit. The function $\text{Var}(\cdot)$ represents the variance. The unit position in the assembly coordinate frame is given by $\mathbf{P}_{i}=[x_{i},y_{i}]$. The desired force and net torque of the MARS are denoted as $F_{assembly}$ and $\mathbf{M}^{\pm}_{assembly}=[M_{x}^{+},M_{x}^{-},M_{y}^{+},M_{y}^{-}]$, respectively. The current torque generated by the thrusters is represented as $\boldsymbol{\tau}^{\pm}=[\tau_{x}^{+},\tau_{x}^{-},\tau_{y}^{+},\tau_{y}^{-}]$, which is computed as:
@@ II-B1 Partial Reallocation
where $k$ is the number of normal units. $\Delta F_{failed}$ and $\Delta\mathbf{M}_{failed}$ denote the loss of thrust and torque due to rotor failure, respectively, and can be calculated using the control efficiency matrix. Then, the angular velocity $\boldsymbol{\omega}_{i}$ can be determined according to the adaptive configuration control method without failure [11].
@@
where $\mathbf{P}_{n_{r}}$ denotes the original control allocation matrix of a multirotor with $n_{r}$ propellers, and $\mathbf{E}_{i}=\text{diag}\left([n^{-1},|\mathbf{P}_{i}+\mu|\cdot J_{assembly}/|\sum_{i}^{n}(\mathbf{P}_{i}+\mu)\cdot J_{unit}|]\right)\in\mathbb{R}^{4\times 4}$ defines the adaptive configuration control allocation parameters. Here, $J_{assembly},J_{unit}$ represent the moment of inertia for MARS and individual units, respectively. $\mu$ is a small constant introduced to prevent division by zero.
@@ II-B2 Full Reallocation
${\rm s.t.}\$$\Sigma f_{i,j}\cdot\mathbf{P}_{i,j}=0$ (6b)
@@ III Collision-free Trajectory Planning
11:    if $N_{near}\notin C$ and CollisionCheck($N_{near}$then
@@ III-1 Discrete sequence generation
Figure 3: (a1)–(e1): Trajectory tracking results with and without FTC. (a2)–(e2): Average position errors across ten repeated trials under different fault scenarios. (a), (b): FTC applied to a $3\times 2$ assembly with one faulty unit. (c), (d), (e): FTC applied to a $3\times 2$ assembly with two faulty units. The faulty propellers are marked red.
@@
where $R(\phi)$, $R(\theta)$, and $R(\psi)$ represent the rotation matrices about the $z$-, $y$-, and $x$-axes, respectively. The maximum force contributions for each unit and the entire assembly are defined as: $\boldsymbol{u}_{\mathcal{M},i}^{\text{max}}=\begin{bmatrix}f_{i1}^{\text{max}},&f_{i2}^{\text{max}},&f_{i3}^{\text{max}},&f_{i4}^{\text{max}}\end{bmatrix}^{T},\quad\boldsymbol{u}_{a}^{\text{max}}=\begin{bmatrix}\boldsymbol{u}_{\mathcal{M},1}^{\text{max}},&\cdots,&\boldsymbol{u}_{\mathcal{M},n}^{\text{max}}\end{bmatrix}^{T}\in\mathbb{R}^{4n}$.
@@ III-2 Dynamical trajectory
Figure 4: FTC performance during self-reconstruction in $3\times 1$ and $3\times 3$ plus assemblies. Faulty propellers are highlighted in red, while orange arrows indicate the range of oscillation. (a) $3\times 1$ assembly (comparison method). (b) $3\times 1$ assembly (Ours). (c) $3\times 3$ plus assembly (comparison method). (d) $3\times 3$ plus assembly (Ours). (e) and (f) Yaw angle and IMU-measured accelerations of the faulty unit during self-reconfiguration FTC in $3\times 1$ and $3\times 3$ plus assemblies, respectively. Note that (a) and (c) are sequences from the original video in [10], which uses the Crazyflie simulator. Moreover, the previous work [10] does not specify the fault-tolerant control method used, and the code is not open-source; therefore, we are unable to evaluate it within the same simulator environment.
@@ IV-A1 Complete Failure of One and Two Units
To evaluate the performance of the proposed method, a spiral trajectory tracking experiment is conducted. Fig. 3 presents the results of trajectory tracking along with the average position errors over 10 repeated trials. As shown in Fig. 3(a) and (b), when fault-tolerant control is enabled, the assembly remains capable of effectively tracking the spiral direction. Although the tracking trajectory deviates from the target due to increased fault-induced torque when faulty units are positioned far from the center, previous works [10, 15] have proposed effective approaches to mitigate these drawbacks. In contrast, without fault-tolerant control, the assembly quickly diverges from the intended trajectory after take-off and fails to complete the tracking task. These results demonstrate that the proposed fault-tolerant control method effectively redistributes power to handle fault scenarios.
Figure 5: Simulation of collision-free trajectory planning for $4\times 1$ and $3\times 2$ assemblies. The annotated diagram illustrates the configuration of MARS, where faulty units are marked in red. The simulation videos are available at https://github.com/RuiHuangNUS/MARS-FTCP/. (a) Normal assembly (SVSDF) [6]. (b) Normal assembly (Ours). (c) Post-failure assembly with FTC (SVSDF). (d) Post-failure assembly with FTC (Ours).
@@
IV-A2 Rotor Failure and Performance Degradation
[…]
@@ IV-C Collision-free Path Planning
[…]
The simulation sequence graphs of different methods are shown in Fig. 5. FTC enables MARS to track the planned trajectory after the failure of a unit in both methods, as illustrated in Fig. 5 (c) and (d). However, the trajectory generated by SVSDF poses greater challenges for MARS with dynamic models, often resulting in more pronounced trajectory errors and a higher frequency of collisions, as shown in Fig. 5 (a) and (c). Specifically, in the first row of Fig. 5, a failure occurs in Unit No. 3 of the $4\times 1$ assembly. Our method demonstrates lower tracking errors compared to [6]. When a failure occurs in Unit No. 4, the increased fault torque reduces the assembly’s control capability, leading to higher tracking errors and even collisions, as shown in the second row of Fig. 5 (c) for the comparison method [6].
Furthermore, for the $3\times 2$ assembly, MARS without fault struggles to accurately follow the SVSDF-generated trajectory, as seen in the third row of Fig. 5 (a). The performance further deteriorates in the presence of faults, as illustrated in Fig. 5 (c) in the third row. In contrast, our approach, shown in Fig. 5 (b) and (d) in the third row, effectively mitigates collision risks and reduces average trajectory errors, achieving performance comparable to the nominal configuration. Table II summarizes the trajectory tracking errors shown in Fig. 5. Across all configurations, our method achieves improvements of at least 48.5% and up to 79.5%, with an average improvement of 69.7%. These results demonstrate the superiority of the proposed method in ensuring dynamic adaptability and collision-free trajectory planning.
TABLE II: Trajectory Error Analysis: Collision-free Path Planning
  Shape (Fault ID) Baseline[6] ($m$) Ours ($m$) Improvement
  $4\times 1$ (Normal) 0.5536 0.2849 $48.5\%$
$4\times 1$ (No. 3) 0.7499 0.3039 $59.5\%$
$4\times 1$ (No. 4) 0.9923 0.3068 $69.1\%$
$3\times 2$ (Normal) 1.1777 0.3216 $72.7\%$
$3\times 2$ (No. 2) 1.7007 0.3486 $79.5\%$
Average 1.0348 0.3132 $69.7\%$
@@ V Conclusion
To achieve robust and safe post-failure control of MARS, this paper proposes a partial and full fault-tolerant power allocation method that redistributes the expected force and torque across all units based on moment arm magnitudes. Compared to existing approaches, our method significantly reduces oscillations during the reconfiguration process. Furthermore, we develop a dynamics-aware, collision-free trajectory planning strategy by extending the A* algorithm to compute optimal attitude nodes that maximize control performance. We then generate a dynamic trajectory incorporating a dynamic feasibility term. Experimental results demonstrate that our approach effectively reduces tracking errors while preventing collisions when following a collision-free trajectory, even in the presence of failures, achieving an average improvement of $69.7\%$.